OpenStudy (anonymous):

I have no idea how to do this What is the equation of the line passing through (8, –6) and perpendicular to the line y = x – 2 in slope-intercept form?

OpenStudy (anonymous):

We know that the slope of our original equation is 1. What we need is the equation of a line that is perpendicular to y = x - 2. The "negative reciprocal" of the slope of our original equation of y=x-2 can be determined by taking the original slope of: \[\frac{ 1 }{ 1}\] and then flipping the numerator and the denominator. The amazing thing is this still gives us 1. However, it is the "negative" reciprocal. So, we multiply by -1. so, the beginning of our new equation is: \[y = -1 x + b\] Then, sub in (8, -6) into your new equation: \[-6=-1(8) + b\] we find that after solving for b: b=2 and our final equation is: y=-x+2

OpenStudy (anonymous):

That was very helpful! Thank you so much!